A military jet first flies in one direction, turns sharply and then flies in another, as described by the vectors: A=i+4j; B=5i-3j. Take the x-axis as east and find the angle in degrees of the sum of these motions relative to east.(use arctan or arccos, no need to calculate).
Answer #1

Net displacement = A + B = (1+45)+ (56-37) 5 = 60 + o = tan (Dy ) o = tan (t) = 9.46 Angle o= 9.46°